Find the present and future value of $1000 received every month end for 20 years if the interest rate is J12 = 12% p.a.
1
Expert's answer
2020-11-01T17:27:50-0500

I) PVOA is present value of an annuity


"=Payment ×\frac{1 - (1+r) ^{-n}} {r}"


="1,000×\frac{1 - (1+0.01) ^{-240}} {0.01}"


PVOA =$90,819.42


ii) FVOA is the future value of the Annuity


"=Payment ×\frac{(1+r) ^{n}-1} {r}"


="Payment ×\frac{(1+0.01) ^{240}-1} {0.01}"


FVOA=$989,255.37